Cost-Based Pricing Decision

Jeremy Costa, owner of Costa Cabinets Inc., is preparing a bid on a job that requires $2,430 of direct materials, $2,187 of direct labor, and $1,458 of overhead. Jeremy normally applies a standard markup based on cost of goods sold to arrive at an initial bid price. He then adjusts the price as necessary in light of other factors (e.g., competitive pressure). Last year’s income statement is as follows:

Sales $195,000
Cost of goods sold 81,900
Gross margin $113,100
Selling and administrative expenses 46,300
Operating income $66,800

Required:

1. Calculate the markup that Jeremy will use. Round your answer to one decimal place.
%

2. What is Jeremy's initial bid price? Round your answer to the nearest dollar.
$
Answer #1

Part 1 ( Markup )

Markup (on sales ) = (Gross Margin / Sales)*100

= (113100 / 195000)*100  

= 58%

Markup (on cost of good sold ) = (Gross Margin / Cost of good sold)*100

= (113100 / 81900)*100  

= 138.1%

Part 2 ( Initial Bid Price )

Cost of Job = Direct Material + Direct Labour + Overhead

= 2430 + 2187 + 1458 = 6075

Bid Price = Cost of Job + Markup on Cost

= 6075 + ( 6075*138.1 )

= 14,465 ( Rounded to nearest dollar )

Initial bid price for the job is $ 14,465